Derive an expression for adiabatic lapse rate.
1
Expert's answer
2019-05-08T09:31:19-0400

Hydrostatic equation

"dp=-g\rho dz"

First law of thermodynamics for adiabatic process

"mc_V dT - \frac{Vdp}{\gamma}=0"

Where

"\gamma=\frac{c_p}{c_V}"

Combining the Hydrostatic equation with first law of thermodynamics for adiabatic process, we have

"c_p dT - \frac{m}{V}dp=0""c_p dT + \frac{m}{V}g\rho dz=0""\frac{dT}{dz}=-\frac{g}{c_p}"

This an expression for adiabatic lapse rate

"\frac{dT}{dz}=-\frac{g}{c_p}"